OpenStudy (anonymous):

Please help! How would I solve t - 2 ( 3 - 2t) = 2t + 9?

OpenStudy (aripotta):

first distribute the -2 into the parentheses

hartnn (hartnn):

use distributive property a(b+c) =ab+ac so whats -2(3-2t) = ?

OpenStudy (anonymous):

-6 - 4t?

hartnn (hartnn):

-2(-2t) = 4t

OpenStudy (anonymous):

so did I do it wrong?

hartnn (hartnn):

so, -2(3-2t) == -6+4t ok?

hartnn (hartnn):

so t-6+4t = 2t+9 can u solve this ?

OpenStudy (anonymous):

would it simplify to -6 + 5t = 2t + 9 first?

hartnn (hartnn):

yes! thas correct. go ahead.

hartnn (hartnn):

add 6 to both sides

OpenStudy (anonymous):

does t = 5?

hartnn (hartnn):

yup! good work

OpenStudy (anonymous):

what about the t in the very front of the equation?

hartnn (hartnn):

which ?